OpenStudy (anonymous):

Is each situation a linear or a nonlinear model? . The stopping distance for a car is given by the equation , where D is the distance a car travels once the brakes are applied and v is the car’s velocity. 2. A home’s value is given by the equation y = 154,000(1.04)x, where y is the value of the house and x is the number of years since the home was purchased.

OpenStudy (solomonzelman):

are you sure that this is the full question (without any information missing) ?

OpenStudy (anonymous):

The first situation is missing a formula. The situation 2 we have: y = 154000*(1.04)^x, right? Think about the graph of this function... is it linear? (i.e. a line)

OpenStudy (anonymous):

If you know enought physics, we can find a equation of the distance D in finction of v: we have from the energy conservation theory that: (m*v^2)/2 = m*g*u*D, where m is mass of the car, g is gravitational acceleration, u is the kinect friction coefficient. D = (v^2)/(2*m*g*u) From that, if we plot this function, will the be a line? (i.e. linear)?

OpenStudy (solomonzelman):

I usually see cases where the velocity function is just a derivative, so I would have all the rights to assume that #1 's function is not linear... or it would usually not be linear

OpenStudy (anonymous):

If you know enought physics, we can find a equation of the distance D in finction of v: we have from the energy conservation theory that: (m*v^2)/2 = m*g*u*D, where m is mass of the car, g is gravitational acceleration, u is the kinect friction coefficient. D = (v^2)/(2*g*u) From that, if we plot this function, will the be a line? (i.e. linear)?

OpenStudy (anonymous):

You are right @SolomonZelman , as we can see: \[m\frac{ dv }{dt } = -mg\mu\]\[\frac{ dv }{ dx }\frac{ dx }{ dt } = -g\mu\]\[v \frac{ dv }{ dx } = -g\mu\]\[\int\limits_{v}^{0}vdv = -\int\limits_{0}^{D}g\mu dx\]\[\frac{ v^2 }{ 2 } = g\mu D\]\[D = \frac{ v^2 }{ 2g\mu }\] As i found above using energy